Please help me I have tried this problem over and over again and this is the answer I came up with.(-2In(c))(In(a)-3In(b))^-3
-3(5)+(2+3)^-3
-15+(6)^-3
-15+-216=-231
Ans.-231


Problem (d) (ln c^−2)(ln a/b^−3)^−3=

Answer by Alan3354(69443)

Please help me I have tried this problem over and over again and this is the answer I came up with.(-2In(c))(In(a)-3In(b))^-3
-3(5)+(2+3)^-3
= -15 + 5^(-3)
= -15 + 1/(5^3)
= -15 + 1/125
= -1874/125
*****************************
-15+(6)^-3
-15+-216=-231
Ans.-231
=====================
Problem (d) (ln c^-2)(ln a/b^-3)^-3
(ln c^-2)(ln a/b^-3)^-3
-----
Confirm that you're multiplying logs & raising logs to exponents.